OpenStudy (missingno):

The area of a rectangle is cube root of 81 feet squared and the length is cube root of 3 feet. What is the width of the rectangle?

OpenStudy (mathmale):

Area = Length times Width. \[area=\sqrt[3]{81};~side~length=\sqrt[3]{3}\]

OpenStudy (mathmale):

Divide the side length into the area to find the width:\[\frac{ \sqrt[3]{81} }{ \sqrt[3]{3} }=?\]

OpenStudy (missingno):

k give me a min

OpenStudy (missingno):

3 to the 4th power?